# SpaceControl: Introducing Test-Time Spatial Control to 3D Generative Modeling

### v1

@@ Results.
Quantitative results are reported in Table 1, while qualitative results are shown in Figure 4. Both Spice-E and Spice-E-T perform well on chairs and tables but struggle to faithfully generate objects that they were not fine-tuned on (Toys4K). SpaceControl significantly outperforms the baselines in all experiments in terms of Chamfer Distance (CD) to the spatial control, while achieving comparable CLIP-I, FID, and P-FID scores. For completeness, we also report scores for the text-conditioned Trellis using the DiT-XL backbone, which is also the base model used in our SpaceControl. Note that for the sake of simplicity in Tab. 1 we only report results of SpaceControl with $\tau_{0}=6$. However, $\tau_{0}$ can be chosen freely by the user, depending on the desired strength of conditioning. For completeness, we report results for different values of $\tau_{0}$ in Tab. 2. We can see that by increasing the value of $\tau_{0}$ and thus strength of the spatial conditioning, we obtain generations which align more closely to the input spatial control.
@@
Table 2: Analysis of $\mathbf{\tau_{0}}$. The evaluation metrics are L2 Chamfer Distance (CD) and Fréchet Inception Distance (FID). CD quantifies alignment with spatial control, while FID assesses realism. CD scores are scaled by $10^{3}$. We show scores for spatial control given as geometric primitives (P) and meshes (M).
@@
5.2 Qualitative results
Besides Figures 1, 4, and 6, we provide additional qualitative results for object editing in the Appendix, visualizing outputs of different methods conditioned on both coarse and detailed input controls. In general, training-based methods struggle to generate objects in specific poses, whereas SpaceControl consistently produces plausible results. For example, other methods generate a cow with two heads (Spice-E and Spice-E-T), an elephant with an eye on its back (Spice-E), or shapes that fail to strictly follow the spatial conditioning or exhibit low quality (Coin3D).
@@ 5.3 Analysis experiments
While existing methods for 3D spatial conditioning do not provide a way to control its strength, our SpaceControl enables flexible interpolation between different levels of adherence. In this section, we evaluate how the parameter $\tau_{0}$ governs the trade-off between fidelity to the spatial control signal and the realism of the generated asset. Quantitative results are reported in Table 2, using the same metrics and datasets as in Table 1. We further present qualitative results in Fig. 3 and in the Appendix, showing how varying the conditioning strength produces different outcomes. Adjusting $\tau_{0}$ allows users to regulate this trade-off according to their preferences, balancing higher shape quality against stronger adherence to the spatial guidance. Additionally, the plot in Figure 3 (left) illustrates this trade-off on Toys4K, indicating that $\tau_{0}\in[4,6]$ generally provides a good compromise between spatial adherence and shape quality.
@@
SpaceControl supports multi-modal control for 3D asset generation by combining spatial guidance via superquadrics with natural language and optional image conditioning. While the model can synthesize assets using only superquadrics and textual prompts, images are particularly useful for maintaining visual consistency during object edits, as shown in Figure 6(a) and in the Appendix. As we only use image prompts in the Appearance Flow Model of Trellis, they primarily affect texture, with only minor influence on geometry. While this capability originates from the pre-trained Trellis, SpaceControl enables its practical use for cross-modal texture transfer, effectively performing style transfer from 2D images to generated 3D shapes.
@@ 6 Discussion and Conclusion
In summary, our approach introduces the first training-free method that by operating directly in the 3D space is able to spatially condition the generation of high quality assets. Through extensive evaluations and a practical interface, we demonstrate both the effectiveness and usability of our method in real-world creative workflows.
Limitations and Future Work. While SpaceControl enables flexible spatial control via a tunable adherence parameter $\tau_{0}$, this parameter is currently selected manually. Although this supports user-driven control over the realism–faithfulness tradeoff, it complicates automated generation of diverse, high-quality assets without per-instance tuning. Additionally, our current formulation enforces a uniform adherence level across the entire object. Future work could explore part-aware control, allowing users to specify which regions should closely follow the input structure and which can deviate more freely to support creative variation.
Reproducibility Statement. Our approach builds on the open-source Trellis model (Xiang et al., 2025), and our experiments use open-source datasets, namely ShapeNet (Chang et al., 2015) and Toys4k (Stojanov et al., 2021). All experiments are fully reproducible, and upon acceptance we will release the code to support replication of our method and results.
@@ Acknowledgments.
Elisabetta Fedele acknowledges support from the ETH AI Center doctoral fellowship, the Swiss National Science Foundation (SNSF) Advanced Grant 216260 (Beyond Frozen Worlds: Capturing Functional 3D Digital Twins from the Real World), and an SNSF Mobility Grant. Francis Engelmann acknowledges support from an SNSF PostDoc mobility fellowship. Or Litany acknowledges support from the Israel Science Foundation (grant 624/25) and the Azrieli Foundation Early Career Faculty Fellowship. This research was also supported in part by an academic gift from Meta. The authors gratefully acknowledge this support.

### v2

@@ Results.
Quantitative results are reported in Tab. 1, while qualitative results are shown in Fig. 5. Both Spice-E and Spice-E-T perform well on chairs and tables but struggle to faithfully generate objects that they were not fine-tuned on (Toys4K). SpaceControl significantly outperforms the baselines in all experiments in terms of Chamfer Distance (CD) to the spatial control, while achieving comparable CLIP-I, FID, and P-FID scores. For completeness, we also report scores for the text-conditioned Trellis using the DiT-XL backbone, which is also the base model used in our SpaceControl. Note that for the sake of simplicity in Tab. 1 we only report results of SpaceControl with $\tau_{0}=6$. However, $\tau_{0}$ can be chosen freely by the user, depending on the desired strength of conditioning. For completeness, we report results for different values of $\tau_{0}$ in Tab. 2. We can see that by increasing the value of $\tau_{0}$ and thus strength of the spatial conditioning, we obtain generations which align more closely to the input spatial control.
@@
Table 2: Analysis of $\tau_{0}$. We evaluate faithfulness to spatial and textual control via Chamfer Distance (CD, scaled by $10^{3}$) and CLIP-I, respectively. Realism is assessed via FID for texture and P-FID for geometry. Results are reported for spatial control provided as geometric primitives (P) and meshes (M).
@@
Qualitative results
are shown in Fig. 1, 5, and 7. Additional results for object editing are in the Appendix, visualizing outputs of different methods conditioned on both coarse and detailed input controls. In general, training-based methods struggle to generate objects in specific poses, whereas SpaceControl consistently produces plausible results. For example, other methods generate a cow with two heads (Spice-E and Spice-E-T), an elephant with an eye on its back (Spice-E), or shapes that fail to strictly follow the spatial conditioning or exhibit low quality (Coin3D).
@@ 5.2 Analysis experiments
While existing methods for 3D spatial conditioning do not provide a way to control its strength, our SpaceControl enables flexible interpolation between different levels of adherence. In this section, we evaluate how the parameter $\tau_{0}$ governs the trade-off between fidelity to the spatial control signal and the realism of the generated asset. Quantitative results are reported in Tab. 2, using the same metrics and datasets as in Tab. 1. We further present qualitative results in Fig. 4 and in the Appendix, showing how varying the conditioning strength produces different outcomes. Adjusting $\tau_{0}$ allows users to regulate this trade-off according to their preferences, balancing higher shape quality against stronger adherence to the spatial guidance. Additionally, the plot in Fig. 4 (left) illustrates this trade-off on Toys4K, indicating that $\tau_{0}\in[4,6]$ generally provides a good compromise between spatial adherence and shape quality.
@@
SpaceControl supports multi-modal control for 3D asset generation by combining spatial guidance with natural language and optional image conditioning. While the model can synthesize assets using geometric conditioning and textual prompts alone, images are particularly useful for maintaining visual consistency during object edits, as shown in Fig. 7(a) and in the Appendix. Moreover, as we only use image prompts in the Appearance Flow Model of Trellis, they primarily affect texture, with only minor influence on the geometry. While this capability originates from the pre-trained Trellis, SpaceControl enables its practical use for cross-modal texture transfer, effectively performing style transfer from 2D images to generated 3D shapes.
@@ 6 Discussion and Conclusion
In summary, SpaceControl introduces the first training-free framework that operates directly in 3D space to enable precise spatial conditioning for high-quality asset generation.Future Work and Potential. The flexibility of SpaceControl provides several promising directions for future exploration. Currently, the adherence parameter $\tau_{0}$ allows users to manually navigate the realism–faithfulness tradeoff. While this enables fine-grained user control, future work could investigate adaptive or learned scheduling for $\tau_{0}$ based on class-specific geometry or sample density. Furthermore, our formulation establishes a foundation for part-aware control; extending the global adherence mechanism to localized regions would allow for varying degrees of geometric fidelity within a single object.Beyond individual assets, our method’s training-free nature makes it naturally extensible to complex 3D scenes. The integration of scene abstraction techniques like SuperDec (Fedele et al., 2025) or Search3D (Takmaz et al., 2025) offers a path toward generating realistic spatial layouts from geometric primitives. Moreover, incorporating functional scene graphs (Zhang et al., 2025) would allow SpaceControl to respect semantic relationships between objects, ultimately enabling the automated generation of “digital cousins” (Dai et al., 2024) for real-world environments.
Reproducibility Statement. Our approach builds on the open-source Trellis model (Xiang et al., 2025), and our experiments use open-source datasets, namely ShapeNet (Chang et al., 2015) and Toys4k (Stojanov et al., 2021). Our code is publicly available on our project page.
@@ Acknowledgments.
Elisabetta Fedele acknowledges support from the ETH AI Center doctoral fellowship, the Swiss National Science Foundation (SNSF) Advanced Grant 216260 (Beyond Frozen Worlds: Capturing Functional 3D Digital Twins from the Real World), and an SNSF Mobility Grant. Francis Engelmann acknowledges support from an SNSF PostDoc mobility fellowship. Or Litany acknowledges support from the Israel Science Foundation (grant 624/25) and the Azrieli Foundation Early Career Faculty Fellowship. This research was also supported in part by an academic gift from NVIDIA and Meta. The authors also acknowledge the support from a SwissAI Grant for Small Projects and an Academic Grant from NVIDIA.